OpenStudy (anonymous):

Of the 27 players trying out for the school basketball team, 8 are more than 6 feet tall and 7 have good aim. What is the probability that the coach would randomly pick a player over 6 feet tall or a player with a good aim? Assume that no players over 6 feet tall have good aim.

OpenStudy (anonymous):

so 8 are more than 6 feet tall and 7 have good aim, but are the rest just average or something?

OpenStudy (anonymous):

cause 12 players are left over

OpenStudy (anonymous):

idk how to do that

OpenStudy (anonymous):

okay what are the options and I can maybe help explain

OpenStudy (anonymous):

7/15 6/15 7/9 5/9

OpenStudy (anonymous):

@aaldia678

OpenStudy (anonymous):

so there are a total of 15 players that are over 6 feet tall or have good aim. Overall, there are 27 players. The question is asking what are the chances the coach will choose a player over 6 feet tall or a player with good aim. Since there are 15 players with good aim or that are 6 feet tall, you put it into a fraction: \[\left(\begin{matrix}15 \\ 27\end{matrix}\right)\]

OpenStudy (anonymous):

this then \[\frac{ 5 }{ 9 }\]reduces down to

OpenStudy (anonymous):

the chances of the coach choosing a player over six feet tall or with good aim is \[\frac{ 5 }{ 9 }\]

OpenStudy (anonymous):

thanks i really need that

OpenStudy (anonymous):

medal? :)